OpenStudy (ilovebmth1234):

A photon has a wavelength of 6.2 meters. Calculate the energy of the photon in joules. (Planck's constant is 6.626 × 10È³Ê joule seconds, the speed of light is 2.998 × 10§ m/s)

OpenStudy (ilovebmth1234):

A. \[3.564*10^{-43}\] B. \[3.204*10^{-26}\] C.\[41.08*10^{-34}\] D. \[13.702*10^{-42}\]

OpenStudy (alexanderphantomhive):

E = hc / λ E = ((6.626*10^-34)*(2.998*10^8)) / 6.2 E = 3.17*10^-26 J

OpenStudy (alexanderphantomhive):

gosh i really dont know im stumped.

OpenStudy (ilovebmth1234):

that's okay i couldn't figure it out either this equation thing confused me

OpenStudy (ilovebmth1234):

im gonna try your answer and if i get it wrong ill let you know:3

OpenStudy (abhisar):

Energy = \(\Large\sf \frac{HC}{\lambda}\)

OpenStudy (ilovebmth1234):

wait scratch that most of your answer was right

OpenStudy (abhisar):

What?

OpenStudy (abhisar):

Plug in the values, Energy = \(\sf \Large \frac{6.626 \times 10^{-34} \times 3 \times 10^8}{6.2}\)

OpenStudy (ilovebmth1234):

i went for answer B witch was closer to @AlexanderPhantomhive's answer she got and i got it right so the answer is B. 3.204x10^-26

OpenStudy (abhisar):

Yes, that's \(\huge \checkmark\)

OpenStudy (ilovebmth1234):

thank you! both of you for your help :3 i wish i could give both of you medals

OpenStudy (abhisar):

You're most welcome c;